OpenStudy (anonymous):

using the ration test find the radiance of convergence for the power series \sum_{n=0}^{\infty} \frac{(n+5)x^n}{2^n+n}

OpenStudy (anonymous):

\[\sum_{0}^{\infty} (n+5)x ^{n} \div 2 ^{n}+n\]

OpenStudy (fibonaccichick666):

I'm sorry, I cannot remember how to do radius of convergence, but this may help: http://tutorial.math.lamar.edu/Classes/CalcII/RatioTest.aspx

OpenStudy (anonymous):

so i use the ration test and end up with \[x/2 \lim_{n \rightarrow \infty} n^2+6n/n^2+6n+5\]

OpenStudy (anonymous):

\[\sum_{n=0}^\infty \frac{(n+5)x^n}{2^n+n}~~?\]

OpenStudy (anonymous):

Oh I didn't even see the Tex code up there...

OpenStudy (fibonaccichick666):

where did the x/2 come from on that? The tex code will help us knderstand yourwriting

OpenStudy (anonymous):

i figured out its 2. if you get an answer k any finite number, time |x-a| the you put 1/k. Where k=1/2 i just forgot to multiply by the reciprocal

OpenStudy (fibonaccichick666):

Did you use the ratio test? I don't see where you may have

OpenStudy (anonymous):

and i didnt realize that with x^n+1/x^n=1 not x

OpenStudy (fibonaccichick666):

ok, you have me flummoxed. I do not understand what you did @SithsAndGiggles it's yours

OpenStudy (anonymous):

Ratio test gives \[\begin{align*}\lim_{n\to\infty}\left|\frac{(n+6)x^{n+1}}{2^{n+1}+n+1}\times\frac{2^n+n}{(n+5)x^n}\right|&=|x|\lim_{n\to\infty}\left|\frac{2^n+n}{2^{n+1}+n+1}\right|\\\\ &=|x|\lim_{n\to\infty}\left|\frac{1+n2^{-n}}{2+(n+1)2^{-n}}\right|\\\\ &=\frac{|x|}{2}\end{align*}\] Between the last two steps, I think I can assume you can tell that \[\lim_{n\to\infty}\frac{n}{2^n}=0\]

OpenStudy (anonymous):

wrong

OpenStudy (fibonaccichick666):

And me

OpenStudy (fibonaccichick666):

you made an error in your first step. You dropped your x

OpenStudy (fibonaccichick666):

\(\frac{x^{n+1}}{x^n} \not = 1\)

OpenStudy (anonymous):

Well maybe I should clarify. I didn't state that the radius is 2. I just showed that, in order for the series to converge, the limit must be less than 1. So we have \(\dfrac{|x|}{2}<1\), which gives \(|x|<2\) and indeed a radius of 2.

OpenStudy (anonymous):

Not wrong, just not finished :P

OpenStudy (anonymous):

then i guess its cause of the two leading n^2 numbeers where the value of the limit is the ratio of the leading coefficients of the same powers. then 1/k which is the value 1/2

OpenStudy (anonymous):

Alright well we all seem to agree the radius is 2, right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Wonderful!

OpenStudy (fibonaccichick666):

lol yea, siths, Just multiple ways of going about it, but @pmkat14 , you should check your work because on an exam you will get the problem wrong

OpenStudy (fibonaccichick666):

even though you arrived at a correct answer